OpenStudy (dls):

10^0.24=? I want to know how is it solved and not the calculated answer,kthx.

OpenStudy (shubhamsrg):

approx kar lo

OpenStudy (shubhamsrg):

:P

OpenStudy (dls):

How do I.

OpenStudy (shubhamsrg):

hmm, 0.24 ko approximately 0.25 maan lo

OpenStudy (shubhamsrg):

10^0.25 = 10^(1/4)

OpenStudy (shubhamsrg):

sqrt(sqrt10)

OpenStudy (shubhamsrg):

sqrt10 = 3.something pi^2 = 10 is a good approximation

OpenStudy (dls):

bhaad me ja ab :P

OpenStudy (dls):

(Y)

OpenStudy (shubhamsrg):

:)

OpenStudy (dls):

:)

OpenStudy (shubhamsrg):

yes, I got a medal! :')

OpenStudy (dls):

You deserved it.

OpenStudy (shubhamsrg):

:') yeah I worked hard! :'))

OpenStudy (shubhamsrg):

for it *

OpenStudy (dls):

\[\LARGE \sqrt{\frac{1.74 \times 10^{-5}}{5}}\]=?

OpenStudy (shubhamsrg):

sqrt(17.4/5) * 10^-3

OpenStudy (shubhamsrg):

sqrt(3.5) * 10^-3 approx

OpenStudy (shubhamsrg):

1.85 approx * 10^-3

OpenStudy (shubhamsrg):

the calculator agrees BI A TCH ! B|

OpenStudy (dls):

\[\LARGE 16 \times 10^{-3}\frac{1 }{70}\]

OpenStudy (dls):

got it :P

OpenStudy (shubhamsrg):



OpenStudy (shubhamsrg):

regained it! B|

OpenStudy (dls):

lo saare medals :P

OpenStudy (shubhamsrg):

:'))

OpenStudy (dls):

mu me lo sabko :P

OpenStudy (shubhamsrg):

:)

OpenStudy (shubhamsrg):